# Automatic Alignment of Discourse Relations of Different Discourse Annotation Frameworks

### v1

@@ Abstract
Existing discourse corpora are annotated based on different frameworks, which show significant dissimilarities in definitions of arguments and relations and structural constraints. Despite surface differences, these frameworks share basic understandings of discourse relations. The relationship between these frameworks has been an open research question, especially the correlation between relation inventories utilized in different frameworks. Better understanding of this question is helpful for integrating discourse theories and enabling interoperability of discourse corpora annotated under different frameworks. However, studies that explore correlations between discourse relation inventories are hindered by different criteria of discourse segmentation, and expert knowledge and manual examination are typically needed. Some semi-automatic methods have been proposed, but they rely on corpora annotated in multiple frameworks in parallel. In this paper, we introduce a fully automatic approach to address the challenges. Specifically, we extend the label-anchored contrastive learning method introduced by Zhang et al. (2022b) to learn label embeddings during a classification task. These embeddings are then utilized to map discourse relations from different frameworks. We show experimental results on RST-DT (Carlson et al., 2001) and PDTB 3.0 (Prasad et al., 2018).
@@ 1.   Introduction
As discourse annotation has high demand on knowledge about discourse, discourse corpora are costly to create. However, these discourse formalisms typically share similar understanding of discourse relations and their essential role in discourse building. Thus, an option to enlarge discourse corpora is to align the existing discourse corpora so that they can be used jointly. This line of work starts as early as Hovy and Maier (1992), but it remains challenging to uncover the alignment of discourse relations used in different annotation frameworks.
@@
As RST-DT and PDTB have an overlapping subsection, the corresponding PDTB-style annotation on the same text is:
@@
where Argument 1 (Arg1) is shown in italics and Argument 2 (Arg2) is in bold. The annotations in brackets represent relation type, which can be implicit, explicit or others, connective, which is identified or inferred by annotators to encode the relation, and sense label, which is delimited with dots, with the first entry showing the sense label at level 1 (L1 sense), the second entry being the sense label at level 2 (L2 sense) and so on.
There are multiple reasons behind the challenges of this task. First, different formalisms have distinctive assumptions about higher-level structures and discourse units. PDTB focuses on the semantic properties of arguments, and argument identification is performed following the Minimality Principle, which means that only those parts that are necessary and minimally required for understanding a relation are annotated (Prasad et al., 2008). In comparison, elementary discourse units (EDUs) in RST are typically clauses. It has been shown repeatedly that segmentation criteria affect the scope of discourse relations and influence the type of relations that can be attached (Demberg et al., 2019; Benamara and Taboada, 2015; Rehbein et al., 2016).
@@
Another major difference between the two frameworks is that RST enforces a tree structure, and all the EDUs and complex discourse units (CDUs) (spans formed by adjacent elementary discourse units and adjacent lower-level spans) should be connected without crossings, while PDTB only focuses on local relations without commitment to any higher-level structure, as exemplified by the three independent annotations shown above. This distinction adds to the difficulty of exploring correlations of relations between the two frameworks, even though the two corpora are built on overlapping texts.
In addition, in RST-DT, an inventory of 78 relations is used, which can be grouped into 16 classes. These relations can be divided into subject matter relations, which are relations whose intended effect is for readers to recognize them, and presentational relations, which are intended to increase some inclination in readers (Mann and Thompson, 1988). For each relation, only one sense label can be attached. In contrast, PDTB adopts a three-level sense hierarchy, and more than one sense label can be annotated for a pair of arguments. As shown in the example, annotation 2 and annotation 3 are annotations for the same argument pair, but different sense labels are assigned. In previous studies that explore the alignment of RST and PDTB discourse relations, these cases typically require manual inspection to determine the closest matching PDTB relation to RST (Demberg et al., 2019). Moreover, PDTB does not take intentional relations into account but focuses on semantic and pragmatic relations.
The combination of these factors makes it challenging to investigate the relationship between discourse relations of different annotation frameworks. Even in empirical studies that make use of corpora annotated in multiple frameworks in parallel, expert knowledge and manual examination are still required. To tackle the challenge caused by differences in discourse segmentation, Demberg et al. (2019) employ the strong nuclearity hypothesis (Marcu, 2000) [^2] A relation that holds between two spans should also hold between the nuclei of the two spans. to facilitate the string matching process of aligning PDTB arguments and RST EDUs. While this method alleviates the limitation of exact string matching of arguments/EDUs, it relies on a corpus annotated with multiple frameworks in parallel. Furthermore, it is conceivable that the relations left out in their analysis because of violating the principle of strong nuclearity hypothesis are not necessarily irrelevant for the goal of enabling joint usage of RST and PDTB.
@@ 3.   Method
$\mathcal{LEQ}=\frac{1}{K}\sum\limits_{i=0}^{K-1}\widetilde{\mathbf{\textit{M}}}_{ii}$ (10)
@@ 7.   Acknowledgments
We thank the anonymous reviewers for insightful feedback and suggestions. Our thanks also go to Mark-Jan Nederhof for fruitful discussion and Craig Myles for his suggestion of using the diagonal entries of the normalized correlation matrix as a metric.

### v2

@@ Abstract
Existing discourse corpora are annotated based on different frameworks, which show significant dissimilarities in definitions of arguments and relations and structural constraints. Despite surface differences, these frameworks share basic understandings of discourse relations. The relationship between these frameworks has been an open research question, especially the correlation between relation inventories utilized in different frameworks. Better understanding of this question is helpful for integrating discourse theories and enabling interoperability of discourse corpora annotated under different frameworks. However, studies that explore correlations between discourse relation inventories are hindered by different criteria of discourse segmentation, and expert knowledge and manual examination are typically needed. Some semi-automatic methods have been proposed, but they rely on corpora annotated in multiple frameworks in parallel. In this paper, we introduce a fully automatic approach to address the challenges. Specifically, we extend the label-anchored contrastive learning method introduced by Zhang et al. (2022b) to learn label embeddings during discourse relation classification. These embeddings are then utilized to map discourse relations from different frameworks. We show experimental results on RST-DT (Carlson et al., 2001) and PDTB 3.0 (Prasad et al., 2018).
@@ 1.   Introduction
As discourse annotation has a high demand on knowledge about discourse, discourse corpora are costly to create. However, these discourse formalisms typically share similar understanding of discourse relations and their role in discourse construction. Thus, an option to enlarge discourse corpora is to align the existing discourse corpora so that they can be used jointly. This line of work starts as early as Hovy and Maier (1992), but it remains challenging to uncover the relationship between discourse relations used in different annotation frameworks.
@@
As RST-DT and PDTB have an overlapping section of annotated texts, the corresponding PDTB-style annotation on the same text is:
@@
where Argument 1 (Arg1) is shown in italics and Argument 2 (Arg2) is in bold. The annotations in parentheses represent relation type, which can be implicit, explicit or others, connective, which is identified or inferred by annotators to signal the relation, and sense label, which is delimited with dots, with the first entry showing the sense label at level 1 (L1 sense), the second entry being the sense label at level 2 (L2 sense) and so on.
The task presents a challenge owing to a multitude of factors. First, different formalisms have distinctive assumptions about higher-level structures and discourse units. PDTB focuses on semantic relations between arguments, and argument identification is performed following the Minimality Principle, which means that only those parts that are necessary and minimally required for understanding a relation are annotated (Prasad et al., 2008). In comparison, elementary discourse units (EDUs) in RST are typically clauses. It has been shown repeatedly that segmentation criteria affect the scope of discourse relations and influence the type of relations that can be attached (Demberg et al., 2019; Benamara and Taboada, 2015; Rehbein et al., 2016).
@@
Another major difference between the two frameworks is that RST enforces a tree structure, and all the EDUs and complex discourse units (CDUs) (spans formed by adjacent elementary discourse units and adjacent lower-level spans) should be connected without crossings, while PDTB only focuses on local relations without commitment to any higher-level structure, as exemplified by the three independent annotations shown above. Previous studies (Lee et al., 2006, 2008) suggest that edge crossings and relations with shared arguments are common for PDTB. This distinction adds to the difficulty of exploring correlations of relations between the two frameworks, even if the two corpora are built on the same texts.
In addition, in RST-DT, an inventory of 78 relations is used, which can be grouped into 16 classes. These relations can be divided into subject matter relations (informational relations in Moore and Pollack (1992)), which are relations whose intended effect is for readers to recognize them, and presentational relations, which are intended to increase some inclination in readers (Mann and Thompson, 1988) (intentional relations in Moore and Pollack (1992)). For each relation, only one sense label can be attached. In contrast, PDTB adopts a three-level sense hierarchy, and more than one sense label can be annotated for a pair of arguments. As shown in the example, annotation 2 and annotation 3 are annotations for the same argument pair, but different sense labels are assigned. In previous studies that explore the alignment of RST and PDTB discourse relations, these cases typically require manual inspection to determine the closest matching PDTB relation to RST (Demberg et al., 2019). Moreover, PDTB does not take intentional relations into account but focuses on semantic and pragmatic relations.
The combination of these factors makes it challenging to investigate the relationship between discourse relations of different annotation frameworks. Even in empirical studies that make use of corpora annotated in multiple frameworks in parallel, expert knowledge and manual examination are still required. To tackle the challenge caused by differences in discourse segmentation, Demberg et al. (2019) employ the strong nuclearity hypothesis (Marcu, 2000) [^2] A relation that holds between two spans should also hold between the nuclei of the two spans. to facilitate the string matching process of aligning PDTB arguments and RST segments. While this method alleviates the limitation of exact string matching of arguments/EDUs, it relies on a corpus annotated with multiple frameworks in parallel. Furthermore, it is conceivable that the relations left out in their analysis because of violating the principle of strong nuclearity hypothesis are not necessarily irrelevant for the goal of enabling joint usage of RST and PDTB.
@@ 3.   Method
[…]
@@ 7.   Acknowledgments
We thank the anonymous reviewers for insightful feedback and suggestions. Our thanks also go to Mark-Jan Nederhof for discussions and Craig Myles for the suggestion of using the diagonal entries of the normalized correlation matrix as a metric.